# Negate or Embrace: On How Misalignment Shapes Multimodal Representation Learning

### v3

@@ C.2 Additional Identification Results
We further analyze the linearity of the identified semantic representations by reporting the $R^{2}$ scores obtained from linear regression applied to the learned features for predicting the true latent variables. As shown in Figure 10 (selection bias) and Figure 11 (perturbation bias), the performance of linear regression closely mirrors that of nonlinear regression reported in Figure 3. This strong correspondence suggests that the relationship between the learned representations and the true latent semantic variables is approximately linear, supporting the conclusion that the identified semantic subspace is nearly linear.
@@
Second, we explore the joint effect of selection and perturbation biases by defining a scenario with selection bias $\mathbb{I}_{\theta}=[8]$ and perturbation bias $\mathbb{I}_{\rho}=[2]$. Results presented in Table 5 demonstrate that when both biases coexist, their effects on semantic identification remain consistent: semantic variables that are either omitted or perturbed are discarded, while unbiased semantic variables—those that are both selected and unperturbed—are reliably preserved in the learned representations.
@@ D.1 Detailed Experimental Setup
MPI3D-Complex (Gondal et al., 2019) contains 460,800 real-world images of resolution $64\times 64\times 3$, spanning all combinations of seven mutually independent, discrete factors (see Table 6). We designate horizontal and vertical positions (hori., vert.) as image-specific due to their low-level visual nature, while treating the remaining five as semantic variables for learning and evaluation purposes.
@@
Text descriptions are generated from the ground-truth semantic attributes using content-word mappings (see Table 6). Under unbiased settings, all five semantic attributes are included. For text-specific variation, a generation template is chosen via a discrete latent variable $m_{t}\sim\mathrm{Uniform}(\{1,2,3\})$, each corresponding to a different human-written sentence structure.
@@
We generate text for each image under various selection and perturbation bias settings to investigate the effects of misalignment. To ensure computational tractability—since exhaustively enumerating all possible configurations is both infeasible and unnecessary—we adopt an incremental strategy for introducing biases into the latent semantic variables, as outlined in Table 7.
@@
For each setting, the dataset is partitioned into training, evaluation, and test subsets in a fixed ratio of $44{,}720:23{,}040:23{,}040$. Across all configurations, we train the image and text encoders for $200{,}000$ steps using the training subset, averaging results over three random seeds. The training objective is the multimodal contrastive loss $\mathcal{L}_{\text{MMCL}}$, defined in Eq. 1, and optimized using the Adam optimizer with an initial learning rate of $1\times 10^{-5}$, a batch size of 256, and a temperature parameter $\tau=1.0$.
@@
For text encoding, we tokenize text using the nltk.PunktTokenizer, following the procedure in (Daunhawer et al., 2022). Tokenized sequences are transformed into two-dimensional one-hot embeddings and processed using a convolutional neural network (CNN) with a variable number of layers, determined by the shape of the tokenized input. The output dimensionality of the CNN is configured to match that of the image encoder, ensuring compatibility in the shared representation space.
@@ Appendix E Experiment Details on Causal3DIdent Dataset
We provide additional details on the Causal3DIdent dataset that are not fully covered in § 5.3. In § E.1, we offer a comprehensive description of the experimental setup, including the image and text latent factors, the image generation process, the design of selection and perturbation bias settings for text generation, and training configurations. In § E.2, we present supplementary results, including analyses of the learned text representations and assessments of the linearity of the learned embeddings.
@@ E.1 Detailed Experimental Setup
Following prior work (Zimmermann et al., 2021; Von Kügelgen et al., 2021; Daunhawer et al., 2022; Yao et al., 2023), we utilize the Causal3DIdent dataset to synthesize images from a predefined latent causal structure. Images are generated using the Blender renderer (Community, 2018), which applies a complex rendering function parameterized by 11 input variables. In our configuration, the object’s $z$-position is fixed, leaving 10 latent factors that govern image generation.
@@
We discretize the continuous variables color, s_color, and b_color using sampled image semantic variables mapped to distinct color palettes: TABLEAU_COLORS for color, CSS4_COLORS for s_color, and XKCD_COLORS for b_color. While s_color remains continuous in the underlying latent representation, we simulate partial information loss for the spotlight position (s_pos) during the mapping to text. As a result, the generated textual descriptions of s_pos do not constitute an information-preserving transformation.
To introduce text-specific variation, we employ five manually designed templates to generate text from the latent factors under each bias setting, adapting the text rendering pipeline from (Daunhawer et al., 2022). A complete list of latent factors and their types is provided in Table 9.
@@
Across all experimental settings, we train the image and text encoders for $100{,}000$ steps on the training subset, using three different random seeds to ensure robustness. The training objective is the multimodal contrastive loss $\mathcal{L}_{\text{MMCL}}$, defined in Eq. 1, and optimized using the Adam optimizer with an initial learning rate of $1\times 10^{-5}$, a batch size of 256, and a temperature parameter $\tau=1.0$. For both the image and text encoders, we adopt the same architectures used in the MPI3D-Complex experiments. The encoding dimensionality is adjusted according to the bias setting: for selection settings 1 through 7, the encoding sizes are set to 3, 3, 4, 5, 5, 6, and 7, respectively; for perturbation settings 1 through 7, the encoding sizes are assigned in reverse order: 7, 6, 5, 5, 4, 3, and 3.
@@
After training the representations, we freeze the encoders and, for each bias setting, train both a linear classifier (logistic regression) and a nonlinear classifier (a two-layer MLP with ReLU activation) for each discrete latent factor. Similarly, for continuous latent factors, we train both a linear regressor and a nonlinear regressor (a two-layer MLP with ReLU activation). All classifiers and regressors are trained for $10{,}000$ steps using the evaluation subset. We assess the predictive performance of the learned representations by evaluating their ability to recover the ground-truth latent factors corresponding to their respective modalities. This evaluation accounts for the fact that some semantic variables may appear in discrete or continuous form, depending on the modality and rendering process. Prediction performance is measured on the test set using the Matthews Correlation Coefficient (MCC) for discrete factors and the coefficient of determination ($R^{2}$) for continuous factors. Metrics are computed separately for each latent factor.
@@ E.2 Additional Results
For other latent semantic variables that are continuous in the image modality but discretized in the text modality—such as color, s_color, and b_color—we observe varying degrees of performance degradation. This drop in performance is likely attributable not only to the quantization of continuous values but also to semantic ambiguity introduced during text generation. Notably, all three attributes correspond to different aspects of color, yet they may be described using overlapping vocabulary drawn from distinct color palettes. For instance, tab:cyan from TABLEAU_COLORS refers to object color (color), cyan from CSS4_COLORS describes spotlight color (s_color), and xkcd:cyan from XKCD_COLORS indicates background color (b_color). Despite referencing different latent variables, these tokens all contain the word cyan, which is tokenized identically by nltk.PunktTokenizer, resulting in ambiguity in the text observations. These findings highlight the importance of using distinct and unambiguous content words when representing semantically different concepts in multimodal learning—particularly when the text modality is not treated merely as an auxiliary input for visual representation learning. Interestingly, the identification of x_pos and y_pos in the text representations does not lead to improved predictability of color, in contrast to what is often observed in the image modality. This aligns with our theoretical expectation that perturbation biases disrupt the underlying causal structure in the image latent space.
Regarding the text-specific factor phrase, we find it to be partially encoded in the learned representations. This contrasts with the image-specific continuous factors, which are consistently omitted. The partial identifiability of phrase is likely attributable to its discrete nature, which violates the conditions typically required for modality-specific factors to be excluded—consistent with findings reported in (Daunhawer et al., 2022). Moreover, this effect appears more pronounced under selection bias settings, particularly when the encoding dimensionality exceeds the true dimensionality of the unbiased semantic subspace.
@@ G.1 Modeling Randomly Missing Latent Semantic Variables
Our current framework addresses misalignment arising primarily from fixed selection and perturbation biases in textual annotations. However, real-world datasets—particularly large-scale, user-generated corpora such as LAION-5B (Schuhmann et al., 2022)—often display random and unstructured semantic omissions. Extending our latent variable model to accommodate such random missingness (e.g., missing completely at random or missing not at random) poses both theoretical and practical challenges. Future research should examine the identifiability consequences of randomly missing semantic variables and establish conditions that ensure partial or probabilistic recovery of latent factors remains feasible.
@@ G.3 Broader Implications
Our findings offer computational support for a contemporary perspective on linguistic relativity (Whorf, 2012), which posits that language shapes human perception and conceptualization. In multimodal AI, linguistic supervision implicitly determines which aspects of the visual world are foregrounded—and which are omitted. Selection and perturbation biases in textual annotations thus act as epistemic filters, constraining the conceptual space that models can represent and generalize over. This reframes dataset design as both an epistemic and normative act: inclusion or omission in annotations encodes a stance on salience, relevance, or appropriateness. Two practical insights follow. First, achieving generalizable representations requires supervision that faithfully captures the intended semantic scope without systematic omissions. Second, curating data to intentionally include or exclude ethically sensitive or socially salient content provides a mechanism for aligning learned representations with human values (Bender and Friedman, 2018; Crawford, 2021). When annotation is treated as an epistemic commitment, dataset design becomes a central tool for shaping both semantic fidelity and ethical alignment.
@@
Beyond explicit annotation content, human annotation errors—such as consistent omissions or mislabels—reveal subtle but powerful behavioral signals. These errors are not uniformly random; rather, they concentrate in dimensions that humans intuitively deprioritize under limited attention (Tversky and Kahneman, 1990; Itti and Koch, 2001). Core attributes like risk or intentionality are rarely mislabeled, while peripheral or low-salience details are more frequently neglected. Such patterns implicitly encode a value hierarchy over semantic factors. This perspective invites a reinterpretation: rather than discarding annotation errors as noise, we can study them as reflections of human value structure. Systematically neglected factors often carry little perceived cost when missed, suggesting lower social or cognitive importance. These regularities form a behavioral prior—informing models not only what remains invariant to environment changes, but what is valued. Incorporating such cues enables a richer form of alignment grounded in human preferences, priorities, and ethical sensibilities (Gabriel, 2020).

### v4

@@ C.2 Additional Identification Results
We further analyze the linearity of the identified semantic representations by reporting the $R^{2}$ scores obtained from linear regression applied to the learned features for predicting the true latent variables. As shown in Figure 10 (selection bias) and Figure 11 (perturbation bias), the performance of linear regression closely mirrors that of nonlinear regression reported in Figure 3. This strong correspondence suggests that the relationship between the learned representations and the true latent semantic variables is approximately linear, indicating that the identified representation subspace is nearly linear.
@@
Second, we explore the joint effect of selection and perturbation biases by defining a scenario with selection bias $\mathbb{I}_{\theta}=[8]$ and perturbation bias $\mathbb{I}_{\rho}=[2]$. Results presented in Table 5 demonstrate that when both biases coexist, their effects on semantic identification remain consistent: semantic variables that are either omitted or perturbed are discarded, while unbiased semantic variables—those that are selected and yet unperturbed—are reliably preserved in the learned representations.
@@ D.1 Detailed Experimental Setup
MPI3D-Complex [18] contains 460,800 real-world images of resolution $64\times 64\times 3$, spanning all combinations of seven mutually independent, discrete factors (see Table 6). We designate horizontal and vertical positions (hori., vert.) as image-specific due to their low-level visual nature, while treating the remaining five as semantic variables for representation learning and evaluation purposes.
@@
Text descriptions are generated from the ground-truth semantic attributes using content-word mappings (see Table 6). Under unbiased settings, all five semantic attributes are included. For text-specific variation, a generation template is chosen via a discrete latent variable $m_{t}\sim\mathrm{Uniform}([3])$, each corresponding to a different human-written sentence structure.
@@
We generate text for each image under various selection and perturbation bias settings to investigate the effects of misalignment. To ensure computational tractability—since exhaustively enumerating all possible configurations is both infeasible and unnecessary—we adopt a progressively incremental strategy for introducing biases into the latent semantic variables, as outlined in Table 7.
@@
For each setting, the dataset is partitioned into training, evaluation, and test subsets in a fixed ratio of $44{,}720:23{,}040:23{,}040$. Across all configurations, we train the image and text encoders for $200{,}000$ steps using the training subset, averaging results over three random seeds. The training objective is the multimodal contrastive loss $\mathcal{L}_{\text{\tiny MMCL}}$, defined in Eq. 1, and optimized using the Adam optimizer with an initial learning rate of $1\times 10^{-5}$, a batch size of 256, and a temperature parameter $\tau=1.0$.
@@
For text encoding, we tokenize text using the nltk.PunktTokenizer, following the procedure in [12]. Tokenized sequences are transformed into two-dimensional one-hot embeddings and processed using a convolutional neural network (CNN) with a variable number of layers, determined by the shape of the tokenized input. The output dimensionality of the CNN is configured to match that of the image encoder, ensuring compatibility in the joint representation space.
@@ Appendix E Experiment Details on Causal3DIdent Dataset
We provide additional details on the Causal3DIdent dataset that are not fully covered in § 5.3. In § E.1, we offer a comprehensive description of the experimental setup, including the image and text latent factors, the image generation process, the design of selection and perturbation bias settings for text generation, and training configurations. In § E.2, we present supplementary results, including analyses of the learned text representations and assessments of the linearity of the learned representations.
@@ E.1 Detailed Experimental Setup
Following prior work [80, 66, 12, 76], we utilize the Causal3DIdent dataset to synthesize images from a predefined latent causal structure. Images are generated using the Blender renderer [8], which applies a complex rendering function parameterized by 11 input variables. In our configuration, the object’s $z$-position is fixed, leaving 10 latent factors that govern image generation.
@@
We discretize the continuous variables color, s_color, and b_color using sampled image semantic variables mapped to distinct color palettes: TABLEAU_COLORS for color, CSS4_COLORS for s_color, and XKCD_COLORS for b_color. While s_color remains continuous in the underlying latent representation, we simulate partial information loss for the spotlight position (s_pos) during the generating mapping to form text. As a result, the generated textual descriptions of s_pos do not constitute an information-preserving transformation.
To introduce text-specific variation, we employ five manually designed templates to generate text from the latent factors under each bias setting, adapting the text rendering pipeline from [12]. A complete list of latent factors and their types is provided in Table 9.
@@
Across all experimental settings, we train the image and text encoders for $100{,}000$ steps on the training subset, using three different random seeds to ensure robustness. The training objective is the multimodal contrastive loss $\mathcal{L}_{\text{\tiny MMCL}}$, defined in Eq. 1, and optimized using the Adam optimizer with an initial learning rate of $1\times 10^{-5}$, a batch size of 256, and a temperature parameter $\tau=1.0$.
For both the image and text encoders, we adopt the same architectures used in the MPI3D-Complex experiments. The encoding dimensionality is adjusted according to the bias setting: for selection settings 1 through 7, the encoding sizes are set to 3, 3, 4, 5, 5, 6, and 7, respectively; for perturbation settings 1 through 7, the encoding sizes are assigned in reverse order: 7, 6, 5, 5, 4, 3, and 3.
@@
After training the representations, we freeze the encoders and, for each bias setting, train both a linear classifier (logistic regression) and a nonlinear classifier (a two-layer MLP with ReLU activation) for each discrete latent factor. Similarly, for continuous latent factors, we train both a linear regressor and a nonlinear regressor (a two-layer MLP with ReLU activation). All classifiers and regressors are trained for $10{,}000$ steps using the evaluation subset.
We assess the predictive performance of the learned representations by evaluating their ability to recover the ground-truth latent factors corresponding to their respective modalities. This evaluation accounts for the fact that some semantic variables may appear in discrete or continuous form, depending on the modality and rendering process.
Prediction performance is measured on the test set using the Matthews Correlation Coefficient (MCC) for discrete factors and the coefficient of determination ($R^{2}$) for continuous factors. Metrics are computed separately for each latent factor.
@@ E.2 Additional Results
For other latent semantic variables that are continuous in the image modality but discretized in the text modality—such as color, s_color, and b_color—we observe varying degrees of performance degradation. This drop in performance is likely attributable not only to the quantization of continuous values but also to semantic ambiguity introduced during text generation. Notably, all three attributes correspond to different aspects of color, yet they may be described using overlapping vocabulary drawn from distinct color palettes. For instance, tab:cyan from TABLEAU_COLORS refers to object color (color), cyan from CSS4_COLORS describes spotlight color (s_color), and xkcd:cyan from XKCD_COLORS indicates background color (b_color). Despite referencing different latent variables, these tokens all contain the word cyan, which is tokenized identically by nltk.PunktTokenizer, resulting in ambiguity in the text observations. These findings highlight the importance of using distinct and unambiguous content words when representing semantically different concepts in multimodal learning—particularly when the text modality is not treated merely as an auxiliary input for visual representation learning.
Interestingly, the identification of x_pos and y_pos in the text representations does not lead to improved predictability of color, in contrast to what is often observed in the image modality. This aligns with our theoretical expectation that perturbation biases disrupt the underlying causal structure in the image latent space.
Regarding the text-specific factor phrase, we find it to be partially encoded in the learned representations. This contrasts with the image-specific continuous factors, which are consistently omitted. The partial identifiability of phrase is likely attributable to its discrete nature, which violates the conditions typically required for modality-specific factors to be excluded—consistent with findings reported in [12]. Moreover, this effect appears more pronounced under selection bias settings, particularly when the encoding dimensionality exceeds the true dimensionality of the unbiased semantic subspace.
@@ G.1 Modeling Randomly Missing Latent Semantic Variables
Our current framework addresses misalignment arising primarily from fixed selection and perturbation biases in textual annotations. However, real-world datasets—particularly large-scale, user-generated corpora such as LAION-5B [59]—often display random and unstructured semantic omissions. Extending our latent variable model to accommodate such random missingness (e.g., missing completely at random or missing not at random) poses both theoretical and practical challenges. Future research should examine the identifiability consequences of randomly missing semantic variables and establish conditions that ensure partial or probabilistic recovery of latent factors remains feasible.
@@ G.3 Broader Implications
Our findings offer computational support for a contemporary perspective on linguistic relativity [68], which posits that language shapes human perception and conceptualization. In multimodal AI, linguistic supervision implicitly determines which aspects of the visual world are foregrounded—and which are omitted. Selection and perturbation biases in textual annotations thus act as epistemic filters, constraining the conceptual space that models can represent and generalize over. This reframes dataset design as both an epistemic and normative act: inclusion or omission in annotations encodes a stance on salience, relevance, or appropriateness. Two practical insights follow. First, achieving generalizable representations requires supervision that faithfully captures the intended semantic scope without systematic omissions. Second, curating data to intentionally include or exclude ethically sensitive or socially salient content provides a mechanism for aligning learned representations with human values [3, 10]. When annotation is treated as an epistemic commitment, dataset design becomes a central tool for shaping both semantic fidelity and ethical alignment.
@@
Beyond explicit annotation content, human annotation errors—such as consistent omissions or mislabels—reveal subtle but powerful behavioral signals. These errors are not uniformly random; rather, they concentrate in dimensions that humans intuitively deprioritize under limited attention [23]. Core attributes like risk or intentionality are rarely mislabeled, while peripheral or low-salience details are more frequently neglected. Such patterns implicitly encode a value hierarchy over semantic factors. This perspective invites a reinterpretation: rather than discarding annotation errors as noise, we can study them as reflections of human value structure. Systematically neglected factors often carry little perceived cost when missed, suggesting lower social or cognitive importance. These regularities form a behavioral prior—informing models not only what remains invariant to environment changes, but what is valued. Incorporating such cues enables a richer form of alignment grounded in human preferences, priorities, and ethical sensibilities [16].